I need to use a linear system to solve this problem;
The perimeter of a rectangular garden is 62 feet. The length is 1 foot more than twice the width. Find the dimension of the garden.

Answer by mananth(16949)

length = x
width = y
2x+2y=62.............1
....
x= 2y+1
x-2y=1.............2
Add (1) & (2)
3x= 63
/3
x=21 the length
x=2y+1
21=2y+1
2y=20
/2
y =10 the width